Find the infinite sum, if it exists: -3 + (0.6) + ( - 0.12)+ .. Preview
Answer #1

The infinite sum formula for GP is:

\frac{a}{1-r}

where a is the first term of GP and r is common ration.

For given list, a= -3 and r= -0.2

So, the infinite sum is:

\frac{-3}{1-(-0.2)}

=-\frac{3}{1.2}

=-2.5

So, the answer is -2.5

Please upvote. Thanks.